# Well-posedness of stochastic mSQG equations with Kraichnan noise and $L^{p}$ data

### v1

@@ 1 Introduction
where $\nabla^{\perp}=(\partial_{2},-\partial_{1})$, $\beta\in(1,2)$, $\circ{\rm d}$ stands for the Stratonovich stochastic differential, and $W=W(t,x)$ is a space-time noise which is white in time, colored and divergence free in space. The existence of weak solutions to (1.1) with $L^{p}$-initial data is relatively classical for suitable choices of $\beta$ and $p>1$. In this paper, we are mainly interested in the uniqueness of solutions to (1.1) in the $L^{p}$-setting. Motivated by Coghi and Maurelli’s work [10] on the stochastic 2D Euler equations, we will show that (1.1) enjoys pathwise uniquenss of weak solutions in suitable spaces, provided that the random perturbation $W(t,x)$ is the famous Kraichnan noise (see [29; 30]), whose covariance matrix $Q(x-y)=\mathbb{E}(W(1,x)\otimes W(1,y))$ is characterized by its Fourier transform:
@@
where $\alpha\in(0,1)$ and $I_{2}$ is the $2\times 2$ identity matrix. We remark that, under very general conditions on $Q$ (see e.g. [21, Section 2.1]), the noise admits the decomposition $W(t,x)=\sum_{k}\sigma_{k}(x)B^{k}_{t}$, where $\{\sigma_{k}\}_{k\geq 1}$ are divergence free vector fields and $\{B^{k}\}_{k\geq 1}$ are independent Brownian motions; thus the first equation in (1.1) can be written more precisely as
@@ 1.1 Motivations
serves as a bridge linking the classical 2D Euler equation in vorticity form ($\beta=2$) and the SQG equation ($\beta=1$), the latter being used in meteorology and oceanography to model the temperature $\theta$ in a rapidly rotating stratified fluid with uniform potential vorticity, cf. [25; 37]. In the influential work [11], Constantin et al. pointed out the structural similarities between the SQG equation and the 3D Euler equation; more precisely, $\nabla^{\perp}\theta$ satisfies an equation which looks similar to the 3D Euler equation in vorticity form. Let us briefly recall a few well-posedness results for the SQG equation, i.e. (1.3) with $\beta=1$. Resnick proved in his PhD thesis [38] the existence of weak solutions to the SQG equation with $L^{2}$-initial data; the result was extended by Marchand [36] to the $L^{p}$-setting with $p>4/3$, leaving open the uniqueness of solutions. Using the techniques of convex integration, Buckmaster et al. [8] proved the nonuniqueness of weak solutions to the SQG equation, even in the presence of fractional dissipation term.
@@
is a fundamental model in fluid dynamics, where $u$ and $\theta$ now represent the fluid velocity and vorticity, respectively. The classical result of Yudovich [43] asserts that if $\theta_{0}\in(L^{1}\cap L^{\infty})(\mathbb{R}^{2})$, then (1.4) admits a unique weak solution in $L^{\infty}([0,T];L^{1}\cap L^{\infty})$. Since then, it has been a challenging open problem in fluid dynamics to show the uniqueness of weak solutions to (1.4) for $(L^{1}\cap L^{p})$-initial data with $p\in(1,\infty)$. In recent years, however, a number of nonuniqueness results for (1.4) appeared. Bressan et al. [6; 7] proposed a (numerically assisted) scheme for showing the existence of nonunique weak solutions to (1.4). By adding a carefully designed forcing term $f$ to (1.4), Vishik [41; 42] constructed nonunique solutions with null initial condition; the construction was revisited by De Lellis and his group in [1] where the authors improved many of the arguments. Albritton et al. [2] adapted the ideas in [1] to construct nonunique Leray-Hopf solutions of 3D Navier-Stokes equations with an external force.
On the other hand, the growing theory of regularization by noise demonstrates that the addition of suitable noises can improve the solution theory of many systems, see [15; 24] for surveys of some early results and the introductions of [10; 21] for more references. In the case of fluid dynamics equations like (1.4), it is by now generally accepted that the transport noise in Stratonovich form is a physically well motivated random perturbation (cf. [20; 26]). Thus, the problem is to search for appropriate space-time noises $W(t,x)=\sum_{k}\sigma_{k}(x)B^{k}_{t}$ such that the solutions to
@@ 1.2 Main results
First, we give the Itô formulation of (1.1) with Kraichnan noise $W(t,x)=\sum_{k}\sigma_{k}(x)B^{k}_{t}$. Thanks to the explicit expression (1.2), one can show that (see [10, Section 2.3])
@@
Let $\alpha\in(0,1)$, $\beta\in(1,2)$ satisfy $2\alpha<\beta<2<\alpha+\beta$, and $\max\big{\{}\frac{2}{1+\beta/2-\alpha},\frac{4}{\beta+1}\big{\}}<p\leq 2$, then for any $\theta_{0}\in\dot{H}^{-\frac{\beta}{2}}\cap L^{p}$, the stochastic mSQG equation (1.7) admits a weak solution in the sense of Definition 1.1, satisfying in addition
@@
    Note that $\beta>2\alpha$ implies $\frac{2}{1+\beta/2-\alpha}<2$, and $\beta>1$ implies $\frac{4}{\beta+1}<2$, so the range of $p$ is not empty. Next, the condition $\alpha+\beta>2$ will be used in Lemma 3.9 below to yield the $L^{2}\big{(}[0,T]\times\Omega,H^{-\frac{\beta}{2}+1-\alpha}\big{)}$-regularity of solutions.
@@
Let $\alpha\in(0,1)$, $\beta\in(1,2)$ satisfy $2\alpha<\beta<2<\alpha+\beta$, and $p>\max\big{\{}\frac{2}{1+\beta/2-\alpha},\frac{3}{\beta}\big{\}}$. Assume $\max\big{\{}0,\frac{2}{p}-\frac{\beta}{2}\big{\}}<\alpha<\frac{\beta}{2}-\frac{1}{p\wedge 2}$ and $\theta_{0}\in L^{1}\cap L^{p}$. Then the following assertions hold.
@@
    $\mathcal{X}:=L^{\infty}\big{(}\Omega\times[0,T],L^{1}\cap L^{p}\big{)}\cap L^{2}\big{(}\Omega\times[0,T],H^{-\frac{\beta}{2}+1-\alpha}\big{)}.$
    More precisely, if $\theta^{1},\,\theta^{2}$ are two weak solutions to (1.7) in the sense of Definition 1.1, on the common filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t},\mathbb{P})$, with the same sequence of Brownian motions $(B^{k})_{k}$ and the same initial condition $\theta_{0}$, then $\mathbb{P}$-a.s. for all $t\in[0,T]$, $\theta^{1}_{t}=\theta^{2}_{t}$.
@@
Recall the Sobolev embedding in dimension 2: $L^{q}\subset\dot{H}^{1-\frac{2}{q}}$ for any $1<q\leq 2$, see e.g. [3, Corollary 1.39]. By the restrictions on $p$, one easily sees that $L^{1}\cap L^{p}\subset\dot{H}^{-\frac{\beta}{2}}$ and the class $\mathcal{X}$ is contained in $L^{\infty}\big{(}[0,T],L^{2}(\Omega,\dot{H}^{-\frac{\beta}{2}})\big{)}$. In the same way, if we additionally assume $p\geq\frac{2}{\beta/2+\alpha}$ in the statements of Theorem 1.4, then the class $L^{2}\big{(}\Omega\times[0,T],H^{-\frac{\beta}{2}+1-\alpha}\big{)}$ is redundant in $\mathcal{X}$.
@@
Fix $p=2$ in Theorem 1.4. Then the restrictions on $\alpha$ and $\beta$ reduce to $\alpha\in(0,1)$, $\beta\in(\frac{3}{2},2)$, $\beta>\max\{2\alpha,2-\alpha\}$ and $1-\frac{\beta}{2}<\alpha<\frac{\beta}{2}-\frac{1}{2}$. Note that $\beta>2-\alpha$ implies $1-\frac{\beta}{2}<\alpha$, so the last two requirements can be combined as $\beta>\max\{2-\alpha,1+2\alpha\}$, which is equivalent to $2-\beta<\alpha<\frac{\beta-1}{2}$. For this purpose, we need $\beta>\frac{5}{3}$.
In summary, for $\beta\in(\frac{5}{3},2)$, we can find a Kraichnan noise $W=\sum_{k}\sigma_{k}B^{k}$ with parameter $\alpha\in(0,\frac{1}{2})$, such that the stochastic mSQG equation (1.7) admits a pathwise unique solution for any initial data $\theta_{0}\in L^{1}\cap L^{2}$. Such a uniqueness result is not known for the deterministic mSQG equation (1.3), thus the addition of noise improves the solution theory for this equation.
@@ 2 Preliminaries
In this section, we present essential prerequisites used in the sequel, including some properties of the covariance of the noise and the kernel related to the nonlinear part in the mSQG equation (1.7). Then we follow the ideas in [10, Section 3] to smooth the irregular elements in (1.7).
@@ 2.1 Mollifying the covariance $Q$
The following properties about the covariance matrix are known, see e.g. [10, Section 2] and [33, Section 10].
@@
$\beta_{N}$ $=(1+\alpha)\beta_{L}>\beta_{L}>0,$
@@
$\begin{split}\sigma_{k}^{\delta}&=\rho^{\delta}*\sigma_{k},\\ Q^{\delta,h}&=\rho^{\delta}*Q,\\ Q^{\delta}&=\rho^{\delta}*\rho^{\delta}*Q.\end{split}$
@@ 2.2 Regularizing the kernel $G_{\beta}$
It is known that $\hat{G}_{\beta}(\cdot)=(2\pi|\cdot|)^{-\beta}$, hence we have $(-\Delta)^{-\frac{\beta}{2}}f=G_{\beta}*f$ for all $f\in\mathcal{S}^{\prime}(\mathbb{R}^{2})$. For more details, see [39, Chapter V, Section 1]. In this paper, we are mainly concerned with the case $21<\beta<$ 2. The derivative and second derivative of $G_{\beta}$ are
@@
$\langle\theta,G_{\beta}^{\delta}*\theta\rangle\rightarrow\int_{\mathbb{R}^{2}}(2\pi|\xi|)^{-\beta}|\hat{\theta}(\xi)|^{2}\,\mathrm{d}\xi=\langle\theta,G_{\beta}*\theta\rangle=(2\pi)^{-\beta}\|\theta\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}.$ (2.10)
@@ 2.3 Smoothing the initial data
The proof of this claim can be found in [10, Section 3].
@@ 3.1 Regularized model
where $c_{\delta}$ is the constant in (2.3). This is a Vlasov-type equation with smooth interaction kernel and smooth common noise. The properties of the solution to equation (3.1) are listed in the next lemma. For a detailed proof, we refer the reader to [10, Appendix C].
@@
$\theta_{t}^{\delta}=\theta_{0}^{\delta}-\int_{0}^{t}\mathrm{div}((K_{\beta}^{\delta}*\theta_{r}^{\delta})\theta_{r}^{\delta})\,\mathrm{d}r-\sum_{k}\int_{0}^{t}\mathrm{div}(\sigma_{k}^{\delta}\theta_{r}^{\delta})\,\mathrm{d}B_{r}^{k}+\frac{c_{\delta}}{2}\int_{0}^{t}\Delta\theta_{r}^{\delta}\,\mathrm{d}r,$
@@
$\mathbb{E}\bigg{[}\sup\limits_{t\in[0,T]}\big{|}\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}*\theta_{t}^{\delta}\rangle\big{|}\bigg{]}\rightarrow 0,\quad\text{ as}\delta\rightarrow 0.$
@@
$\begin{split}&\Big{|}\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}*\theta_{t}^{\delta}\rangle\Big{|}\\ &=\int_{\mathbb{R}^{2}}|\xi|^{-\beta}(1-e^{-(2\pi|\xi|)^{\beta}\delta})|\widehat{\theta_{t}^{\delta}}(\xi)|^{2}\,\mathrm{d}\xi+\int_{\mathbb{R}^{2}}|\xi|^{-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}|\widehat{\theta_{t}^{\delta}}(\xi)|^{2}\,\mathrm{d}\xi\\ &\leq(2\pi)^{\beta}\delta\int_{\mathbb{R}^{2}}\lvert\widehat{\theta_{t}^{\delta}}(\xi)\rvert^{2}\,\mathrm{d}\xi+\int_{\mathbb{R}^{2}}|\xi|^{-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}|\widehat{\theta_{t}^{\delta}}(\xi)|^{2}\,\mathrm{d}\xi.\end{split}$
@@
$|\widehat{\theta_{t}^{\delta}}(\xi)|\leq|\widehat{\theta_{t}^{\delta}}(0)|+\|\nabla\widehat{\theta_{t}^{\delta}}\|_{L^{\infty}}|\xi|\leq 2\pi|\xi|\int_{\mathbb{R}^{2}}|x|\,|\theta_{t}^{\delta}(x)|\,\mathrm{d}x.$
@@
$\int_{\mathbb{R}^{2}}|\xi|^{-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}|\widehat{\theta_{t}^{\delta}}(\xi)|^{2}\,\mathrm{d}\xi$ $\leq 4\pi^{2}\int|\xi|^{2-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}\bigg{(}\int|x|\,|\theta_{t}^{\delta}(x)|\,\mathrm{d}x\bigg{)}^{2}\,\mathrm{d}\xi$
@@
$\left|\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}*\theta_{t}^{\delta}\rangle\right|\lesssim\delta\bigg{(}\|\theta_{t}^{\delta}\|_{L^{2}}^{2}+\Big{(}\int|x|\,|\theta_{t}^{\delta}(x)|\,\mathrm{d}x\Big{)}^{2}\bigg{)}.$
By Lemma 3.1, $\|\theta_{t}^{\delta}\|_{L^{2}}^{2}\leq\|\theta_{0}^{\delta}\|_{L^{2}}^{2}\leq\|\theta_{0}^{\delta}\|_{L^{1}}\|\theta_{0}^{\delta}\|_{L^{\infty}}$; by [10, Lemma C.1], we get
@@
$\mathbb{E}\bigg{[}\sup\limits_{t\in[0,T]}\left|\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}*\theta_{t}^{\delta}\rangle\right|\bigg{]}$
@@
$\sup\limits_{t\in[0,T]}\big{\|}|\nabla(G_{\beta}-G_{\beta}^{\delta})|*|\theta_{t}^{\delta}|\big{\|}_{L^{\infty}}\lesssim\delta^{\frac{\beta-1}{\beta+3}}(\|\theta_{0}^{\delta}\|_{L^{\infty}}+\|\theta_{0}^{\delta}\|_{L^{1}}).$
@@
$|\nabla(G_{\beta}-G_{\beta}^{\delta})|*|\theta_{t}^{\delta}|(x)$ $=\int_{\mathbb{R}^{2}}|\nabla(G_{\beta}-G_{\beta}^{\delta})(x-y)|\,|\theta_{t}^{\delta}(y)|\,\mathrm{d}y$

### v2

@@ 1 Introduction
where $\nabla^{\perp}=(\partial_{2},-\partial_{1})$, $\beta\in(1,2)$, $\circ{\rm d}$ stands for the Stratonovich stochastic differential, and $W=W(t,x)$ is a space-time noise which is white in time, colored and divergence free in space. The existence of weak solutions to (1.1) with $L^{p}$-initial data is relatively classical for suitable choices of $\beta$ and $p>1$. In this paper, we are mainly interested in the uniqueness of solutions to (1.1) in the $L^{p}$-setting. Motivated by Coghi and Maurelli’s work [11] on the stochastic 2D Euler equations, we will show that (1.1) enjoys pathwise uniquenss of weak solutions in suitable spaces, provided that the random perturbation $W(t,x)$ is the famous Kraichnan noise (see [30; 31]), whose covariance matrix $Q(x-y)=\mathbb{E}(W(1,x)\otimes W(1,y))$ is characterized by its Fourier transform:
@@
where $\alpha\in(0,1)$ and $I_{2}$ is the $2\times 2$ identity matrix. We remark that, under very general conditions on $Q$ (see e.g. [22, Section 2.1]), the noise admits the decomposition $W(t,x)=\sum_{k}\sigma_{k}(x)B^{k}_{t}$, where $\{\sigma_{k}\}_{k\geq 1}$ are divergence free vector fields and $\{B^{k}\}_{k\geq 1}$ are independent Brownian motions; thus the first equation in (1.1) can be written more precisely as
@@ 1.1 Motivations
serves as a bridge linking the classical 2D Euler equation in vorticity form ($\beta=2$) and the SQG equation ($\beta=1$), the latter being used in meteorology and oceanography to model the temperature $\theta$ in a rapidly rotating stratified fluid with uniform potential vorticity, cf. [26; 38]. In the influential work [12], Constantin et al. pointed out the structural similarities between the SQG equation and the 3D Euler equation; more precisely, $\nabla^{\perp}\theta$ satisfies an equation which looks similar to the 3D Euler equation in vorticity form. Let us briefly recall a few well-posedness results for the SQG equation, i.e. (1.3) with $\beta=1$. Resnick proved in his PhD thesis [39] the existence of weak solutions to the SQG equation with $L^{2}$-initial data; the result was extended by Marchand [37] to the $L^{p}$-setting with $p>4/3$, leaving open the uniqueness of solutions. Using the techniques of convex integration, Buckmaster et al. [9] proved the nonuniqueness of weak solutions to the SQG equation, even in the presence of fractional dissipation term.
@@
is a fundamental model in fluid dynamics, where $u$ and $\theta$ now represent the fluid velocity and vorticity, respectively. The classical result of Yudovich [44] asserts that if $\theta_{0}\in(L^{1}\cap L^{\infty})(\mathbb{R}^{2})$, then (1.4) admits a unique weak solution in $L^{\infty}([0,T];L^{1}\cap L^{\infty})$. Since then, it has been a challenging open problem in fluid dynamics to show the uniqueness of weak solutions to (1.4) for $(L^{1}\cap L^{p})$-initial data with $p\in(1,\infty)$. In recent years, however, a number of nonuniqueness results for (1.4) appeared. Bressan et al. [7; 8] proposed a (numerically assisted) scheme for showing the existence of nonunique weak solutions to (1.4). By adding a carefully designed forcing term $f$ to (1.4), Vishik [42; 43] constructed nonunique solutions with null initial condition; the construction was revisited by De Lellis and his group in [1] where the authors improved many of the arguments. Albritton et al. [2] adapted the ideas in [1] to construct nonunique Leray-Hopf solutions of 3D Navier-Stokes equations with an external force.
On the other hand, the growing theory of regularization by noise demonstrates that the addition of suitable noises can improve the solution theory of many systems, see [16; 25] for surveys of some early results and the introductions of [11; 22] for more references. In the case of fluid dynamics equations like (1.4), it is by now generally accepted that the transport noise in Stratonovich form is a physically well motivated random perturbation (cf. [21; 27]). Thus, the problem is to search for appropriate space-time noises $W(t,x)=\sum_{k}\sigma_{k}(x)B^{k}_{t}$ such that the solutions to
@@ 1.2 Main results
First, we give the Itô formulation of (1.1) with Kraichnan noise $W(t,x)=\sum_{k}\sigma_{k}(x)B^{k}_{t}$. Thanks to the explicit expression (1.2), one can show that (see [11, Section 2.3])
@@
Let $\alpha\in(0,1)$, $\beta\in(1,2)$ satisfy $1-\frac{\beta}{2}<\alpha<\frac{\beta}{2}$, and $\max\big{\{}\frac{2}{1+\beta/2-\alpha},\frac{4}{\beta+1}\big{\}}<p\leq 2$, then for any $\theta_{0}\in\dot{H}^{-\frac{\beta}{2}}\cap L^{p}$, the stochastic mSQG equation (1.7) admits a weak solution in the sense of Definition 1.1, satisfying in addition
@@
    Note that $\beta>2\alpha$ implies $\frac{2}{1+\beta/2-\alpha}<2$, and $\beta>1$ implies $\frac{4}{\beta+1}<2$, so the range of $p$ is not empty. Next, the condition $22\alpha+\beta>$ 2 will be used in Lemma 3.9 below to yield the $L^{2}\big{(}[0,T]\times\Omega,H^{-\frac{\beta}{2}+1-\alpha}\big{)}$-regularity of solutions.
@@
Let $\alpha\in(0,\frac{1}{2})$, $\beta\in(\frac{3}{2},2)$ satisfy $\frac{2}{p\wedge 2}-\frac{\beta}{2}<\alpha<\frac{\beta}{2}-\frac{1}{p\wedge 2}$ for some $p>\frac{3}{\beta}$. Assume $\theta_{0}\in L^{1}\cap L^{p}$, then the following assertions hold.
@@
    $\mathcal{X}:=L^{\infty}\big{(}\Omega\times[0,T],L^{1}\cap L^{p}\big{)}.$
    More precisely, if $\theta^{1},\,\theta^{2}\in\mathcal{X}$ are two weak solutions to (1.7) in the sense of Definition 1.1, with the same initial condition $\theta_{0}$ and the same sequence of Brownian motions $(B^{k})_{k}$ on the common filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t},\mathbb{P})$, then $\mathbb{P}$-a.s. for all $t\in[0,T]$, $\theta^{1}_{t}=\theta^{2}_{t}$.
@@
The conditions $\alpha\in(0,\frac{1}{2})$ and $\beta\in(\frac{3}{2},2)$ imply $\frac{3}{\beta}>\frac{4}{\beta+1}>\frac{2}{1+\beta/2-\alpha}$. Recall the Sobolev embedding in dimension 2: $L^{q}\subset\dot{H}^{1-\frac{2}{q}}$ for any $1<q\leq 2$, see e.g. [4, Corollary 1.39]. By the restriction $p>\frac{3}{\beta}$, one easily sees that $L^{1}\cap L^{p}\subset\dot{H}^{-\frac{\beta}{2}}$ and the class $\mathcal{X}$ is contained in $L^{\infty}\big{(}[0,T],L^{2}(\Omega,\dot{H}^{-\frac{\beta}{2}})\big{)}$, hence all conditions in Theorem 1.2 are satisfied here. In the same way, the condition $\alpha>\frac{2}{p}-\frac{\beta}{2}$, or equivalently $p>\frac{2}{\beta/2+\alpha}$, implies that the class $\mathcal{X}$ is contained in $L^{2}\big{(}\Omega\times[0,T],H^{-\frac{\beta}{2}+1-\alpha}\big{)}$.
@@
The results in Theorem 1.4 can be understood as follows: for every fixed $\beta\in(\frac{3}{2},2)$, assume that the initial datum $\theta_{0}$ belongs to the space $L^{1}\cap L^{p}$ with $p>\frac{3}{\beta}$, then we can find a special multiplicative transport noise, namely the Kraichnan noise with some suitable index $\alpha$, such that the 2D stochastic mSQG equation (1.7) admits a pathwise unique solution. Such a uniqueness result is not known for the deterministic mSQG equation (1.3), thus the addition of noise improves the solution theory for this equation.
Almost at the same time of our work, Bagnara, Galeati and Maurelli [3] studied the same regularization problem. They assume $p\geq 2$ and are able to treat the full range of $\beta\in(1,2)$ (note that $\beta$ here is related to the parameter in their paper by the relation $\beta\mapsto 2-\beta$); for $\beta$ close to 1, the parameter $p$ has to be big enough.
@@ 2 Preliminaries
In this section, we present essential prerequisites used in the sequel, including some properties of the covariance of the noise and the kernel related to the nonlinear part in the mSQG equation (1.7). Then we follow the ideas in [11, Section 3] to smooth the irregular elements in (1.7).
@@ 2.1 Mollifying the covariance $Q$
The following properties about the covariance matrix are known, see e.g. [11, Section 2] and [34, Section 10].
@@
$\beta_{N}$ $=(1+2\alpha)\beta_{L}>\beta_{L}>0,$
@@
$\begin{split}\sigma_{k}^{\delta}&=\rho^{\delta}\ast\sigma_{k},\\ Q^{\delta,h}&=\rho^{\delta}\ast Q,\\ Q^{\delta}&=\rho^{\delta}*\rho^{\delta}\ast Q.\end{split}$
@@ 2.2 Regularizing the kernel $G_{\beta}$
It is known that $\hat{G}_{\beta}(\cdot)=(2\pi|\cdot|)^{-\beta}$, hence we have $(-\Delta)^{-\frac{\beta}{2}}f=G_{\beta}*f$ for all $f\in\mathcal{S}^{\prime}(\mathbb{R}^{2})$. For more details, see [40, Chapter V, Section 1]. In this paper, we are mainly concerned with the case $21<\beta<$ 2. The derivative and second derivative of $G_{\beta}$ are
@@
$\langle\theta,G_{\beta}^{\delta}\ast\theta\rangle\rightarrow\int_{\mathbb{R}^{2}}(2\pi|\xi|)^{-\beta}|\hat{\theta}(\xi)|^{2}\,\mathrm{d}\xi=\langle\theta,G_{\beta}\ast\theta\rangle=(2\pi)^{-\beta}\|\theta\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}.$ (2.10)
@@ 2.3 Smoothing the initial data
The proof of this claim can be found in [11, Section 3].
@@ 3.1 Regularized model
where $c_{\delta}$ is the constant in (2.3). This is a Vlasov-type equation with smooth interaction kernel and smooth common noise. The properties of the solution to equation (3.1) are listed in the next lemma. For a detailed proof, we refer the reader to [11, Appendix C].
@@
$\theta_{t}^{\delta}=\theta_{0}^{\delta}-\int_{0}^{t}\mathrm{div}((K_{\beta}^{\delta}\ast\theta_{r}^{\delta})\theta_{r}^{\delta})\,\mathrm{d}r-\sum_{k}\int_{0}^{t}\mathrm{div}(\sigma_{k}^{\delta}\theta_{r}^{\delta})\,\mathrm{d}B_{r}^{k}+\frac{c_{\delta}}{2}\int_{0}^{t}\Delta\theta_{r}^{\delta}\,\mathrm{d}r,$
@@
$\mathbb{E}\bigg{[}\sup\limits_{t\in[0,T]}\Big{|}\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}\ast\theta_{t}^{\delta}\rangle\Big{|}\bigg{]}\rightarrow 0,\quad\text{as }\delta\rightarrow 0.$
@@
$\begin{split}&\Big{|}\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}\ast\theta_{t}^{\delta}\rangle\Big{|}\\ &=\int_{\mathbb{R}^{2}}|\xi|^{-\beta}\big{(}1-e^{-(2\pi|\xi|)^{\beta}\delta}\big{)}\big{|}\widehat{\theta_{t}^{\delta}}(\xi)\big{|}^{2}\,\mathrm{d}\xi+\int_{\mathbb{R}^{2}}|\xi|^{-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}\big{|}\widehat{\theta_{t}^{\delta}}(\xi)\big{|}^{2}\,\mathrm{d}\xi\\ &\leq(2\pi)^{\beta}\delta\int_{\mathbb{R}^{2}}\big{|}\widehat{\theta_{t}^{\delta}}(\xi)\big{|}^{2}\,\mathrm{d}\xi+\int_{\mathbb{R}^{2}}|\xi|^{-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}\big{|}\widehat{\theta_{t}^{\delta}}(\xi)\big{|}^{2}\,\mathrm{d}\xi.\end{split}$
@@
$\big{|}\widehat{\theta_{t}^{\delta}}(\xi)\big{|}\leq\big{|}\widehat{\theta_{t}^{\delta}}(0)\big{|}+\big{\|}\nabla\widehat{\theta_{t}^{\delta}}\big{\|}_{L^{\infty}}|\xi|\leq 2\pi|\xi|\int_{\mathbb{R}^{2}}|x|\,|\theta_{t}^{\delta}(x)|\,\mathrm{d}x.$
@@
$\int_{\mathbb{R}^{2}}|\xi|^{-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}\big{|}\widehat{\theta_{t}^{\delta}}(\xi)\big{|}^{2}\,\mathrm{d}\xi$ $\leq 4\pi^{2}\int|\xi|^{2-\beta}e^{-(2\pi|\xi|)^{\beta}/\delta}\bigg{(}\int|x|\,|\theta_{t}^{\delta}(x)|\,\mathrm{d}x\bigg{)}^{2}\,\mathrm{d}\xi$
@@
$\left|\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}\ast\theta_{t}^{\delta}\rangle\right|\lesssim\delta\bigg{(}\|\theta_{t}^{\delta}\|_{L^{2}}^{2}+\Big{(}\int|x|\,|\theta_{t}^{\delta}(x)|\,\mathrm{d}x\Big{)}^{2}\bigg{)}.$
By Lemma 3.1, $\|\theta_{t}^{\delta}\|_{L^{2}}^{2}\leq\|\theta_{0}^{\delta}\|_{L^{2}}^{2}\leq\|\theta_{0}^{\delta}\|_{L^{1}}\|\theta_{0}^{\delta}\|_{L^{\infty}}$; by [11, Lemma C.1], we get
@@
$\mathbb{E}\bigg{[}\sup\limits_{t\in[0,T]}\left|\|\theta_{t}^{\delta}\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}-(2\pi)^{\beta}\langle\theta_{t}^{\delta},G_{\beta}^{\delta}\ast\theta_{t}^{\delta}\rangle\right|\bigg{]}$
@@
$\sup\limits_{t\in[0,T]}\big{\|}|\nabla(G_{\beta}-G_{\beta}^{\delta})|\ast|\theta_{t}^{\delta}|\big{\|}_{L^{\infty}}\lesssim\delta^{\frac{\beta-1}{\beta+3}}(\|\theta_{0}^{\delta}\|_{L^{\infty}}+\|\theta_{0}^{\delta}\|_{L^{1}}).$
@@
$|\nabla(G_{\beta}-G_{\beta}^{\delta})|\ast|\theta_{t}^{\delta}|(x)$ $=\int_{\mathbb{R}^{2}}|\nabla(G_{\beta}-G_{\beta}^{\delta})(x-y)|\,|\theta_{t}^{\delta}(y)|\,\mathrm{d}y$